1.You are performing a left-tailed z-test

If α=.005α=.005, and your test statistic is z=−2.59, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis

You are performing a left-tailed z-test

If α=.10, and your test statistic isz=-2.01, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis

You are performing a right-tailed t-test with a sample size of 15
If α=.01, and your test statistic is t=3.02t=3.02, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis

You are performing a right-tailed z-test

If α=.005, and your test statistic is z=1.96, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis

You are performing a left-tailed z-test

If α=.10, and your test statistic is z=-1.48, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis
Answer #1

Solution! Given that, d=0.005 a test statisticis z=-2.59 This is left-tailed test - P(Z <-2.59) = 0.0048 P-value = 0.0048 0.0n=15 L=o.ol dr=n-1=15-1 = 14 - t =3.02 This is right-tailed test p-value = 1 -(+3.02;14) -|-0.9954 -0.0046 p-value =0.00 46 0d= o.10 Z=-1.48 This is left tailed test PC25-1•48) = 0.0694 p-value = 0.0694 0.0694 Lo.10 p-value ad Reject Ho Reject Ho nul